A 43.2 sample of CaCO3 was treated with an excess of aqueous H2SO4, producing calcium sulfate (CaSO4), 3.65 g of water and CO2(g). What was the % yield of H2O?
1
Expert's answer
2019-11-20T06:54:24-0500

According to the reaction:

CaCO3 + H2SO4 = CaSO4 + H2O + CO2

From here, the theoretical yield of water equals::

mt (H2O) = m(CaCO3) × Mr(H2O) / Mr(CaCO3) = 43.2 g × 16 g/mol / 100 g/mol = 6.912 g

As the actual (ma) yield of water equals 3.65 g, the percent yield is:

% yield = (ma / mt) × 100% = (3.65 g / 6.912 g) × 100% = 52.81%


Answer: 52.81%